# Rethinking Efficient and Effective Point-based Networks for Event Camera Classification and Regression: EventMamba

### v2

@@ Abstract
Event cameras, drawing inspiration from biological systems, efficiently detect changes in ambient light with low latency and high dynamic range while consuming minimal power. The most current approach to processing Event Cloud often involves converting it into frame-based representations, which neglects the sparsity of events, loses fine-grained temporal information, and increases the computational burden. In contrast, Point Cloud is a popular representation for 3D processing and serves as an alternative method to address the limitations mentioned above. Nevertheless, previous point-based methods show an unsatisfactory performance compared to the frame-based method. In order to bridge the gap, we propose EventMamba, an efficient and effective framework based on Point Cloud representation that achieves competitive results even compared to the state-of-the-art (SOTA) frame-based method in both classification and regression tasks. This notable accomplishment is facilitated by our rethinking of the distinction between Event Cloud and Point Cloud, emphasizing effective temporal information extraction through optimized network structures. Specifically, EventMamba leverages temporal aggregation and State Space Model (SSM) based Mamba boasting enhanced temporal information extraction capabilities. Through a hierarchical structure, EventMamba is adept at abstracting local and global spatial features and implicit and explicit temporal features. By adhering to the lightweight design principle, EventMamba delivers impressive results with minimal computational resource utilization, demonstrating its efficiency and effectiveness. In the experiments, EventMamba exhibits SOTA point-based performance on five different scales of action recognition datasets and distinguishes itself on both Camera Pose Relocalization (CPR) and eye-tracking regression tasks, outperforming all frame-based methods.
@@ 1 Introduction
Event cameras present an attractive option for deploying lightweight networks capable of real-time applications. Unlike traditional cameras, which sample light intensity at fixed intervals, event pixels independently respond to in-situ gradient light intensity changes, indicating the corresponding point’s movement in three-dimensional space [1]. This unique capability allows event cameras to capture sparse and asynchronous events that respond rapidly to changes in brightness. Event cameras demonstrate remarkable sensitivity to rapid light intensity changes, detecting log intensity in individual pixels and producing event outputs typically on the order of a million events per second (eps). In comparison, standard cameras can only capture 60 frames per second, underscoring the exceptional speed of event cameras [2]. Furthermore, the distinctive operating principle of event cameras grants them superior power efficiency, low latency, less redundant information, and high dynamic range. Consequently, event cameras have found diverse applications in high-speed target tracking, simultaneous localization and mapping (SLAM), as well as in industrial automation.
@@
Event Cloud is a data structure that encapsulates the coordinates, timestamps, and polarity of each event occurrence [3]. The temporal stream can discern fast behaviors with microsecond resolution [4]. Several studies have strived to augment and refine the problem of processing dynamic motion information from event cameras. Figure 2 (a) illustrates the frame-based method, which partitions the Event Cloud into small chunks and accumulates the events for each time interval onto the image plane [5]. The grayscale on the plane denotes the intensity of the event occurrence at that coordinate. However, the frame-based method disregards the fine-grained temporal information within a frame and may generate blurred images in fast-moving action scenarios. Moreover, this approach escalates the data density and computes even in event-free regions, forfeiting the sparse data benefits of an event-based camera. Consequently, the frame-based method is unsuitable for scenarios with fast-moving actions and cannot be implemented with limited computational resources.
@@
Figure 1: The gain in three datasets is visualized. EventMamba achieves SOTA in the point-based method and maintains very high efficiency. The shapes illustrate the results corresponding to different data formats, with varying colors denoting Floating Point Operations Per Second (FLOPs).
@@
In order to overcome the issues discussed above, alternative approaches, such as the point-based method have been proposed [6, 7, 8]. Point Cloud is a collection of 3D points $(x,y,z)$ that represents the shape and surface of an object or environment and is often used in lidar and depth cameras [9]. Event Cloud is a collection of events $(x,y,t,p)$ generated by event cameras, $t$ represents timestamps and $p$ is the polarity. By treating each event’s temporal information as the third dimension, event inputs $(x,y,t)$ can be regarded as points and aggregated into a pseudo-Point Cloud [6, 10, 11]. This method is well-suited for analyzing Event Cloud, which can then be directly fed into the Point Cloud network, as shown in Figure 2 (b). By doing so, the data format transformation process is greatly simplified, and the fine-grained temporal information and sparse property are preserved [6].
@@
On the other hand, our primary goal is to democratize computer vision technology by making it accessible to a wider range of devices and applications. It is indeed accurate that the conventional frame-based method generally displays negligible processor power consumption. However, with the advent of DVS technology, the power consumption of sensors has been drastically reduced by 90%-95% [12], rendering it much lower than that of conventional ANN processors. As a result, decreasing the algorithm’s power consumption has the potential to significantly reduce overall power consumption and enable the deployment of numerous battery-powered or resource-constrained devices, such as mobile phones, drones, and IoT devices. To achieve this goal, we introduce a novel Point Cloud framework, named EventMamba, for event-based action recognition [7], CPR, and eye-tracking tasks, whose gains are shown in Figure 1. EventMamba is characterized by its lightweight parameterization and computational demands, making it ideal for resource-constrained devices. It features preprocessing, a hierarchy structure, and a classifier or regressor. First, the Event Cloud is divided into many sliding windows. The classification task is sampled through the subwindows [7], while the regression task is straightforward in accumulating enough points for processing [8]. Then, the sampled points are fed into a hierarchy structure to abstract high-level features. The hierarchy structure comprises local spatial, global spatial, and explicit temporal extractors. The feature extractors are interconnected through temporal aggregation. Finally, the general Multi-Layer Perceptron (MLP) is used for classification and regression. The main contributions of our work are:
1. We directly process the raw data obtained from the event cameras, preserving the sparsity and fine-grained temporal information.
2. We strictly arrange the events in chronological order, which can satisfy the prerequisites to extract the temporal features absent in vanilla Point Cloud Network.
3. We employ temporal aggregation and Mamba modules to capture explicit temporal features during the hierarchy structure effectively.
4. We propose a lightweight framework with a generalization ability to handle different-size event-based classification and regression tasks.
@@ 2.1 Frame-based Method
Frame-based methods have been extensively employed for event data processing. These methods convert events into a density map by integrating the number of events within a specific time interval, which can be visualized as a 2D image [13]. The image can be directly fed into traditional feature extraction techniques as shown in Figure 2 (a) for benchmarking purposes. The following will offer a comprehensive review of frame-based methods applied to three distinct tasks.
@@ 2.1.2 Camera Pose Relocalization
CPR task is an emerging application in power-constrained devices and has gained significant attention. It aims to train several scene-specific neural networks to accurately relocalize the camera pose within the original scene used for training. RGB-based CPR networks have been extensively studied [19]. PoseNet [20] introduced transfer learning by replacing the classifier of GoogLeNet [21] with a regressor specifically designed for absolute pose regression. Building upon this end-to-end concept, researchers have further improved the PoseNet framework by integrating advanced backbone architectures like VGG [22] and ResNet [23]. Alternative strategies have also been explored, such as incorporating LSTM networks as a substitute for the localizer or meticulously designing customized loss functions [24, 25, 26]. Additionally, the Auxiliary Learning methods have significantly contributed to boosting the network’s performance [27, 28, 29]. However, this end-to-end structure often suffers from overfitting, which can adversely impact performance. Recently, hybrid approaches have emerged, combining learning-based techniques for local problems with traditional image retrieval or structure-based pipelines. DSAC series, for instance, achieve high pose estimation accuracy [30, 31], but come with increased computational costs and latency, especially for edge devices.
Event-based CPR methods often derive from the frame-based CPR network. SP-LSTM [32] employed the stacked spatial LSTM networks to process event images, facilitating a real-time pose estimator. To address the inherent noise in event images, Jin Yu et al.[33] proposed a network structure combining denoise networks, convolutional neural networks, and LSTM, achieving good performance under complex working conditions. In contrast to the aforementioned methods, a novel representation named Reversed Window Entropy Image (RWEI) [34] is introduced, which is based on the widely used event surface [35] and serves as the input to an attention-based DSAC* pipeline [30] to achieve SOTA results. However, the computationally demanding architecture involving representation transformation and hybrid pipeline poses challenges for real-time execution. Additionally, all existing methods ignore the fine-grained temporal feature of the event cameras, and accumulate events into frames for processing, resulting in unsatisfactory performance.
@@ 2.1.3 Eye-tracking
Event-based eye-tracking methods leverage the inherent attribution of event data to realize low-latency tracking. Ryan et al. [36] propose a fully convolutional recurrent YOLO architecture to detect and track faces and eyes. Stoffregen et al. [37] design Coded Differential Lighting method in corneal glint detection for eye tracking and gaze estimation. Angelopoulos et al. [38] propose a hybrid frame-event-based near-eye gaze tracking system, which can offer update rates beyond 10kHz tracking. Zhao et al. [39] also leverage the near-eye grayscale images with event data for a hybrid eye-tracking method. The U-Net architecture is utilized for eye segmentation and a post-process is designed for pupil tracking. Chen et al. [40] presents a sparse Change-Based Convolutional Long Short-Term Memory model to achieve high tracking speed in the synthetic event dataset.
@@ 2.2 Point-based Method
With the introduction of PointNet [10], Point Cloud can be handled directly as input, marking a significant evolution in processing technologies. This development is further enhanced by PointNet++ [11], which integrates the Set Abstraction (SA) module. This module serves as an innovative tool to elucidate the hierarchical structure related to global and local Point Cloud processing. Although the original PointNet++ relies on basic MLPs for feature extraction, there have been considerable advancements to enhance Point Cloud processing methodologies [42, 43, 44]. Furthermore, an intriguing observation is that even simplistic methodologies can yield remarkable outcomes. A case in point is PointMLP [45], despite its reliance on an elementary ResNet block and affine module, consistently achieves SOTA results in the domains of point classification and segmentation.
@@ 3.1 Event Cloud Preprocessing
The length of the Event Cloud in the dataset exhibits variability. To ensure the consistency of inputs to the network, we employ a sliding window to segment the Event Cloud. In the task of action recognition, it is imperative to include a complete action within a sliding window. Thus, we choose to use the second level for sliding window length. Meanwhile, for the task of Camera Pose Relocalization and eye tracking, the resolution of the pose ground truth is available at the millisecond level. As a result, the length of the sliding window is set to the millisecond level. The splitting process can be represented by the following equation:
@@ 3.2.1 $LocalFE$ Extractor
This extractor comprises a cascade of sampling, grouping, and residual blocks. Aligned with the frame-based design concept, our focus is to capture both local and global information. Local information is acquired by leveraging Farthest Point Sampling (FPS) to sample and K-nearest neighbors (KNN) to the group. The following equations can represent the operations of sampling and grouping.
@@
Subsequently, the residual block is utilized to extract higher-level features. Concatenating residual blocks can indeed enhance the model’s capacity for learning, yet when dealing with sparsely sampled event-based data, striking the right balance between parameters, computational demands, and accuracy becomes a pivotal challenge. In contrast to frame-based approaches, the strength of Point Cloud processing lies in its efficiency, as it necessitates a smaller dataset and capitalizes on the unique event characteristics. However, if we were to design a point-based network with parameters and computational demands similar to frame-based methods, we would inadvertently forfeit the inherent advantages of leveraging Point Clouds for event processing. The Residual block can be written as a formula:
@@ 3.2.2 Temporal Aggregation
$\mathcal{SA}=A\cdot F_{local}=S_{t1}\cdot a_{t1}+S_{t2}\cdot a_{t2}+\dots+S_{tk}\cdot a_{tk},$ (17)
@@ 3.2.3 $GlobalFE$ Extractor
Unlike previous work[8], employing Bi-LSTM for explicit temporal information extraction nearly doubles the number of parameters, despite achieving amazing results. We incorporate the explicit temporal feature extraction module into the hierarchical structure, opting for the faster and stronger Mamba block instead [49]. The topology of the block is shown in the green background of Figure 3. This block mainly integrates the sixth version of the State Space Model (SSM), which is well able to parallel and focus on long time series of information, such as temporal correlations between 512 event groups. The SSM extracts explicit temporal features that can be expressed as follows:
@@
Where $x_{t},h_{t},$ and $y_{t}$ are the SSM’s discrete inputs, states, and outputs. $\mathbf{A}$ and $\mathbf{B}$ are the continuous system parameters, while $\bar{\mathbf{A}}$, $\bar{\mathbf{B}}$ and $\Delta$ are the parameters in the discrete system by the zero-order hold rule. $T^{\prime}$ and $D^{\prime}$ are the number and dimension of events for the current stage, respectively. The whole $GlobalFE$ Extractor can be represented by this formula:
$\mathcal{ST}=\text{Mamba}(\mathcal{S})$ (22)
$\mathcal{ST}=\text{ResB}(\mathcal{ST})$ (23)
@@
3.3 Regreesor and Loss Function
@@
where $\hat{x}$ and $\hat{y}$ stand for the prediction value along the $x$ and $y$ dimensions. $x$ and $y$ are the ground truth value along the x and y dimensions. $w_{x}$ and $w_{y}$ are the hyperparameters to adjust the weight of the x dimension and y dimension in the loss function. $i$ means the $i_{th}$ pixel of the prediction or ground truth label.

### v4

@@ Abstract
Event cameras draw inspiration from biological systems, boasting low latency and high dynamic range while consuming minimal power. The most current approach to processing Event Cloud often involves converting it into frame-based representations, which neglects the sparsity of events, loses fine-grained temporal information, and increases the computational burden. In contrast, Point Cloud is a popular representation for processing 3-dimensional data and serves as an alternative method to exploit local and global spatial features. Nevertheless, previous point-based methods show an unsatisfactory performance compared to the frame-based method in dealing with spatio-temporal event streams. In order to bridge the gap, we propose EventMamba, an efficient and effective framework based on Point Cloud representation by rethinking the distinction between Event Cloud and Point Cloud, emphasizing vital temporal information. The Event Cloud is subsequently fed into a hierarchical structure with staged modules to process both implicit and explicit temporal features. Specifically, we redesign the global extractor to enhance explicit temporal extraction among a long sequence of events with temporal aggregation and State Space Model (SSM) based Mamba. Our model consumes minimal computational resources in the experiments and still exhibits SOTA point-based performance on six different scales of action recognition datasets. It even outperformed all frame-based methods on both Camera Pose Relocalization (CPR) and eye-tracking regression tasks. Our code is available at: https://github.com/rhwxmx/EventMamba.
@@ 1 Introduction
Event cameras present an attractive option for deploying lightweight networks capable of real-time applications. Unlike traditional cameras, which sample light intensity at fixed intervals, event pixels independently respond to in-situ gradient light intensity changes, indicating the corresponding point’s movement in three-dimensional space [1]. This unique capability allows event cameras to capture sparse and asynchronous events that respond rapidly to changes in brightness. Event cameras demonstrate remarkable sensitivity to rapid light intensity changes, detecting log intensity in individual pixels and producing event outputs typically on the order of a million events per second (eps). In comparison, standard cameras can only capture 120 frames per second, underscoring the exceptional speed of event cameras [2]. Furthermore, the distinctive operating principle of event cameras grants them superior power efficiency, low latency, less redundant information, and high dynamic range. Consequently, event cameras have found diverse applications in high-speed target tracking, simultaneous localization and mapping (SLAM), as well as in industrial automation.
@@
The event camera’s output contains coordinates, timestamps, and polarity of each event occurrence [3], which is defined as Event Cloud $(x,y,t,p)$. The temporal stream can discern fast behaviors with microsecond resolution [4]. Several studies have strived to augment and refine the problem of processing dynamic motion information from event cameras. Figure 2 (a) illustrates the frame-based method, which partitions the Event Cloud into small chunks and accumulates the events for each time interval onto the image plane [5, 6]. The grayscale on the plane denotes the intensity of the event occurrence at that coordinate. However, the frame-based method disregards the fine-grained temporal information within a frame and may generate blurred images in fast-moving action scenarios. Moreover, this approach escalates the data density and computes even in event-free regions, forfeiting the sparse data benefits of an event-based camera. Consequently, the frame-based method is unsuitable for scenarios with fast-moving actions and cannot be implemented with limited computational resources.
@@
Figure 1: The gain in three datasets is visualized. EventMamba achieves SOTA in the point-based method and maintains very high efficiency. The shapes illustrate the results corresponding to different data formats, with varying colors denoting Floating Point Operations (FLOPs).
@@
In order to overcome the issues discussed above, alternative approaches, such as the point-based method have been proposed [7, 8, 9]. Point Cloud is a collection of 3D points $(x,y,z)$ that represents the shape and surface of an object or environment and is often used in lidar and depth cameras [10]. By treating each event’s temporal information as the third dimension, event inputs $(x,y,t)$ can be regarded as points and aggregated into a pseudo-Point Cloud [7, 11, 12]. This method is well-suited for analyzing Event Cloud, which can then be directly fed into the Point Cloud network, as shown in Figure 2 (b). By doing so, the data format transformation process is greatly simplified, and the fine-grained temporal information and sparse property are preserved [7].
@@
In this paper, we introduce a novel Point Cloud framework, named EventMamba, for event-based action recognition, CPR, and eye-tracking tasks, whose gains are shown in Figure 1. Consistency with our previous works[8, 9], EventMamba is characterized by its lightweight parameterization and computational demands, making it ideal for resource-constrained devices. Diverging from our previous works, TTPOINT [8], which did not consider temporal feature extraction, and PEPNet [9], which employed A-Bi-LSTM behind the hierarchical structure, we try to integrate the explicit temporal feature extraction module into the hierarchical structure to enhance temporal features, forming a unified, efficient, and effective framework. We also embrace the hardware-friendly Mamba Block, which demonstrates proficiency in processing sequential data. The main contributions of our work are:
  • We directly process the raw data obtained from the event cameras, preserving the sparsity and fine-grained temporal information.
  • We strictly arrange the events in chronological order, which can satisfy the prerequisites to extract the temporal features absent in the vanilla Point Cloud Network.
  • We employ temporal aggregation and Mamba block to enhance explicit temporal features during the hierarchy structure effectively.
  • We propose a lightweight framework with a generalization ability to handle different-size event-based classification and regression tasks.
@@ 2.1 Frame-based Method
Frame-based methods have been extensively employed for event data processing. The following will offer a comprehensive review of frame-based methods applied to three distinct tasks.
@@ 2.1.2 Camera Pose Relocalization
CPR task is an emerging application in power-constrained devices and has gained significant attention. It aims to train several scene-specific neural networks to accurately relocalize the camera pose within the original scene used for training [19]. Event-based CPR methods often derive from the frame-based CPR network [20, 21, 22, 21, 23, 22, 24]. SP-LSTM [25] employed the stacked spatial LSTM networks to process event images, facilitating a real-time pose estimator. To address the inherent noise in event images, Jin Yu et al.[26] proposed a network structure combining denoise networks, convolutional neural networks, and LSTM, achieving good performance under complex working conditions. In contrast to the aforementioned methods, a novel representation named Reversed Window Entropy Image (RWEI) [27] is introduced, which is based on the widely used event surface [28] and serves as the input to an attention-based DSAC* pipeline [24] to achieve SOTA results. However, the computationally demanding architecture involving representation transformation and hybrid pipeline poses challenges for real-time execution. Additionally, all existing methods ignore the fine-grained temporal feature of the event cameras, and accumulate events into frames for processing, resulting in unsatisfactory performance.
@@ 2.1.3 Eye-tracking
Event-based eye-tracking methods leverage the inherent attribution of event data to realize low-latency tracking. Ryan et al. [29] propose a fully convolutional recurrent YOLO architecture to detect and track faces and eyes. Stoffregen et al. [30] design Coded Differential Lighting method in corneal glint detection for eye tracking and gaze estimation. Angelopoulos et al. [31] propose a hybrid frame-event-based near-eye gaze tracking system, which can offer update rates beyond 10kHz tracking. Zhao et al. [32] also leverage the near-eye grayscale images with event data for a hybrid eye-tracking method. The U-Net architecture is utilized for eye segmentation, and a post-process is designed for pupil tracking. Chen et al. [33] presents a sparse Change-Based Convolutional Long Short-Term Memory model to achieve high tracking speed in the synthetic event dataset.
@@ 2.2 Point-based Method
With the introduction of PointNet [11], Point Cloud can be handled directly as input, marking a significant evolution in processing technologies. This development is further enhanced by PointNet++ [12], which integrates the Set Abstraction (SA) module. This module serves as an innovative tool to elucidate the hierarchical structure related to global and local Point Cloud processing. Although the original PointNet++ relies on basic MLPs for feature extraction, there have been considerable advancements to enhance Point Cloud processing methodologies [35, 36, 37, 38, 39]. Furthermore, an intriguing observation is that even simplistic methodologies can yield remarkable outcomes. A case in point is PointMLP [40], despite its reliance on an elementary ResNet block and affine module, consistently achieves SOTA results in the domains of point classification and segmentation.
@@
2.3 State Space Models
State Space Models (SSM) have emerged as a versatile framework initially rooted in control theory, now widely applied due to its superior capability process sequence modeling tasks [44, 45, 46]. SSM-based Mamba [47] has shown promise as a hardware-friendly backbone that can achieve linear-time inference and can be an alternative to the transformer. While primarily focused on language and speech data, recent advancements like S4ND [48], Vision Mamba [49], and PointMamba [37] have extended SSMs to visual tasks, showing potential in competing with established models like ViTs. Nevertheless, neither the image nor Point Cloud contains strict temporal information, requiring both to employ customized serialization and ordering methods to align with the SSM sequences [50]. Event-based works [34, 51] also focus on frame-based representations, which do not take full advantage of the rich and fine-grained temporal information in events. In this paper, we leverage point cloud representations, preserving the temporal sequence of event occurrences, and integrated Mamba with a hierarchical structure to comprehensively extract temporal features between events.
@@ 3.1 Event Cloud Preprocessing
The length of the Event Cloud in the dataset exhibits variability. To ensure the consistency of inputs to the network, we employ a sliding window to segment the Event Cloud. In the task of action recognition, it is imperative to include a complete action within a sliding window. Therefore, we choose to set the sliding window length in seconds. Meanwhile, for the task of Camera Pose Relocalization and eye tracking, the resolution of the pose ground truth is available at the millisecond level. As a result, we chose to set the sliding window length to milliseconds. The splitting process can be represented by the following equation:
@@ 3.2.1 $LocalFE$ Extractor
This extractor comprises a cascade of sampling, grouping, and residual blocks. Aligned with the frame-based design concept [52], our focus is to capture both local and global information. Local information is acquired by leveraging Farthest Point Sampling (FPS) to sample and K-nearest neighbors (KNN) to the group. The following equations can represent the operations of sampling and grouping.
@@
Subsequently, the residual block is utilized to extract higher-level features. Concatenating residual blocks can indeed enhance the model’s capacity for learning, yet when dealing with sparsely sampled event-based data, striking the right balance between parameters, computational demands, and accuracy becomes a pivotal challenge. In contrast to frame-based approaches, the strength of Point Cloud processing lies in its efficiency, as it necessitates a smaller dataset and capitalizes on the unique event characteristics. However, if we were to design a point-based network with parameters and computational demands similar to frame-based methods, we would inadvertently forfeit the inherent advantages of leveraging Point Clouds for event processing. The Residual block (ResB) can be written as a formula:
@@ 3.2.2 Temporal Aggregation
$\mathcal{SA}=\mathcal{S}\cdot A=S_{t1}\cdot a_{t1}+S_{t2}\cdot a_{t2}+\dots+S_{tk}\cdot a_{tk},$ (17)
@@ 3.2.3 $GlobalFE$ Extractor
Unlike previous work[9], employing Bi-LSTM for explicit temporal information extraction nearly doubles the number of parameters, despite achieving amazing results. We incorporate the explicit temporal feature extraction module into the hierarchical structure to enhance temporal features, opting for the faster and stronger Mamba block instead [47]. The topology of the block is shown in the green background of Figure 3. This block mainly integrates the sixth version of the State Space Model (SSM), which is well able to parallel and focus on long time series of information, such as temporal correlations between 1024 event groups. The SSM extracts explicit temporal features that can be expressed as follows:
@@
Where $x_{t},h_{t},$ and $y_{t}$ are the SSM’s discrete inputs, states, and outputs. $\mathbf{A}$, $\mathbf{B}$ and $\mathbf{C}$ are the continuous system parameters, while $\bar{\mathbf{A}}$, $\bar{\mathbf{B}}$ and $\Delta$ are the parameters in the discrete system by the zero-order hold rule. $T^{\prime}$ and $D^{\prime}$ are the number and dimension of events for the current stage, respectively. The whole $GlobalFE$ Extractor can be represented by this formula:
$\mathcal{ST}=$ $\quad\text{Mamba}(\mathcal{SA})$ (22)
$\mathcal{ST}=$ $\quad\text{ResB}(\mathcal{ST})$ (23)
@@
3.3 Regressor and Loss Function
@@
where $\hat{x}$ and $\hat{y}$ stand for the prediction value along the $x$ and $y$ dimensions. $x$ and $y$ are the ground truth values along the x and y dimensions. $w_{x}$ and $w_{y}$ are the hyperparameters to adjust the weight of the x dimension and y dimension in the loss function. $i$ means the $i_{th}$ pixel of the prediction or ground truth label.